Louise spikes a volleyball. At the instant the ball leaves her hand, its height is 2.6 m and its resultant velocity is 20 m/s downward and forward at an angle of 60° below horizontal.

How long will it take for the ball to strike the floor if the opposing team does not block it?

How far will the ball travel horizontally before it strikes the floor?

Make sure to show all work.
Answer #1

downward speed of volleyball =20*sin(60) = 17.32 m/s

horizontal speed of volleyball =20*cos(60) = 10 m/s

2.6 = 17.32*t + 10*t^2

10*t^2 +17.32*t - 2.6 = 0

t = 0.138 seconds

time =  0.138 seconds

distance =  0.138*10 = 1.38 meters